Consider the following data: Two solutions of an unknown slightly soluble salt, A(OH)2, were allowed to equilibrate--one at 25 degrees Celsius and the other at 80 degrees Celsius. A 15.00 mL aliquot of each solution is titrated with 0.250 M HCl. 6.49 mL of the acid is required to reach the endpoint of the titration at 25 degrees Celsius, while 63.07 mL are required for the 80 C solution.

1. Assuming that the change in enthalpy is negligible over this temperature range, calculate deltaH.

2. Assuming that the change in entropy is negligible over this temperature range, calculate deltaS.
Answer #1

Step 1: Calculate Ksp for the reaction at both the given temperatures:

                                               A(OH)2 (s)\rightleftharpoons 2 OH-(aq) + A +(aq)

concentration at equilibrium                                          2x                   x

and Ksp = (2x)2 (x)   = 4x3

At 25 °C , mmoles of HCl used = 0.250 M x 63.07 mL =15.77 = mmoles of OH- present.

total volume (in ml) = 15 + 63.07 = 78.07

[OH-] = 15.77/78.07 = 0.202 M   = 2x , \Rightarrow x = 0.101 M

\RightarrowKsp = 2.19 x 10-4

At 80 °C , mmoles of HCl used = 0.250 M x 6.49 mL =1.62 = mmoles of OH- present.

total volume (in ml) = 15 + 6.49 = 21.49

[OH-] = 1.62/21.49 = 0.075 M   = 2x , \Rightarrow x = 0.038 M

\RightarrowKsp = 4.12 x 10-3

Step 2: Calculate \Delta G at both temperatures, using the formula:

\Delta G = -RT ln (K)

At 25 °C, T= 298 K and R = 8.314 J/mol. K

\Delta G = -(8.314 J/mol. K )x 298 K ln ( 2.19 x 10-4) = 20. 9 kJ/mol

Similarly,

At 80 °C, T= 353 K and R = 8.314 J/mol. K

\Delta G = -(8.314 J/mol. K )x 353 K ln ( 4.12 x 10-3) = 16.12 kJ/mol

Step 3: Using Gibbs Helmholtz equation, that is:

             \frac{\mathrm{d(\frac{\Delta G}{T}) } }{\mathrm{d} T} = -\frac{\Delta H}{T^{2}}

And, Assuming that \Delta H is independent of Temperature in the given range, we can simplify the above equation as:

d\frac{\Delta G}{T} =- \frac{\Delta H}{T^{2}} dT

and integrating the above from T1 to T2,

\frac{\Delta G_{2}}{T_{2}} - \frac{\Delta G_{1}}{T_{1}} = \Delta H (\frac{1}{T_{2}} -\frac{1}{T_{1}})

Putting the values of \Delta G obtained above for T1 = 298K and T2 = 353 K, we have

\frac{16.12}{353} - \frac{20.9}{298} = \Delta H (\frac{1}{353} -\frac{1}{298})

which gives, \Delta H = 46.80 kJ/mol

Step 4: To calculate entropy change,

Use the equation;

\Delta G ( T_{2})= \Delta H - T_{2}\Delta S               ...... Note that H and S are assumed constant for the given Temperature range.

\Delta G ( T_{1})= \Delta H - T_{1}\Delta S

Subtract the second equation from the first one,

\Delta G ( T_{2}) -\Delta G ( T_{1})= T_{1}\Delta S-T_{2}\Delta S

or,

\frac{\Delta G ( T_{2}) -\Delta G ( T_{1})}{(T_{1}-T_{2})}= \Delta S

Putting the values in above equation,

\Delta S = \frac{16.12-20.9}{(298-353)}

\Rightarrow \Delta S = 86.9 J/K

KINDLY SUBMIT FEEDBACK IF EXPLANATION IS CLEAR